### v1

@@ Abstract
In scientific research and its application, scientific literature analysis is crucial as it allows researchers to build on the work of others. However, the fast growth of scientific knowledge has led to a massive increase in scholarly articles, making in-depth literature analysis increasingly challenging and time-consuming. The emergence of Large Language Models (LLMs) has offered a new way to address this challenge. Known for their strong abilities in summarizing texts, LLMs are seen as a potential tool to improve the analysis of scientific literature. However, existing LLMs have their own limits. Scientific literature often includes a wide range of multimodal elements, such as molecular structure, tables, and charts, which are hard for text-focused LLMs to understand and analyze. This issue points to the urgent need for new solutions that can fully understand and analyze multimodal content in scientific literature. To answer this demand, we present Uni-SMART (Universal Science Multimodal Analysis and Research Transformer), an innovative model designed for in-depth understanding of multimodal scientific literature. Through rigorous quantitative evaluation across several domains, Uni-SMART demonstrates superior performance over leading text-focused LLMs. Furthermore, our exploration extends to practical applications, including patent infringement detection and nuanced analysis of charts. These applications not only highlight Uni-SMART’s adaptability but also its potential to revolutionize how we interact with scientific literature.
@@ 1 Introduction
To rigorously assess the multimodal capabilities of Uni-SMART, a comparative analysis was conducted against leading LLMs, such as Gemini, GPT-3.5, and GPT-4. Our assessment targets several data types critical to the comprehension of scientific documents: tables, charts, molecular structures, and chemical reactions. The results demonstrate Uni-SMART’s superior performance in all tested areas, especially in understanding and analyzing complex multimodal contents, thus highlighting its potential as a helpful assistant for scientific literature analysis.
In the following sections, Section 2 details Uni-SMART’s data sources and iterative training approach. Subsequently, Section 3 – Evaluation, presents detailed comparisons of Uni-SMART with leading LLMs across a variety of modalities, showcasing its advanced capabilities in multimodal data interpretation. Then, Section 4 – Application showcases some specific applications of Uni-SMART in the scientific domain. Finally, Section 5 discusses Uni-SMART’s limitations and future research avenues, along with its potential impact on scientific research and technological advancement.
@@ 2 Method
  • Multimodal Learning: During the initialization phase, the model is trained with a limited set of multimodal data to recognize and extract diverse information elements from scientific literature, outputting these in a sequential format that incorporates both textual and multimodal data.
  • LLM SFT: The output sequences, along with corresponding question-answer pairs, are utilized for supervised fine-tuning (SFT) of LLM, boosting its proficiency in handling multimodal content.
  • User Feedback: The SFT-enhanced LLM is deployed in real applications, facilitating the collection of user feedback. Samples receiving positive feedback are subsequently filtered and incorporated into the data enhancement, while those with negative feedback are subject to expert annotation before being integrated into the data enhancement process.
  • Expert Annotation: Samples with negative feedback are meticulously annotated by human experts. This step ensures that the model learns from its mistakes, with semi-automated tools assisting in this process to enhance efficiency. Negative feedback is examined to identify challenges in multimodal recognition or the language model’s reasoning abilities. This analysis helps determine which multimodal elements and types of question are more challenging for the model, facilitating focused improvements.
  • Data Enhancement: Finally, the annotated data, along with partial samples that received positive feedback, are added to the training dataset for data enhancement. This iterative process progressively improves Uni-SMART’s performance, enriching its training dataset and enhancing its multimodal understanding capabilities.
Such cyclical pipeline significantly enhances Uni-SMART’s performance in a variety of challenging tasks, such as information extraction, complex element identification, scientific literature understanding/analysis, and multimodal understanding/reasoning.
@@ 3 Evaluation
In this section, we perform a detailed quantitative evaluation on the capabilities of Uni-SMART and leading LLMs across various modalities. Table 1 presents the statistics of the benchmarks. For details of the specific evaluation tasks, please refer to SciAssess [10]. And the experimental setup for each model is detailed in Appendix A.
Table 1: Dataset statistics.
Modality
Domain
Task
Question #
Table
Alloy Materials
Composition Extraction
55
Drug Discovery
Affinity Data Extraction
15
Organic Materials
Electrolyte Solubility data extraction
8
Electrolyte Table QA
48
Polymer Property Extraction
15
OLED Property Extraction
13
Chart
Alloy Materials
Alloy ChartQA
15
Biology
Biology ChartQA
15
Drug Discovery
Drug ChartQA
15
Organic Materials
Polymer ChartQA
15
Molecule
Drug Discovery
Affinity Data Extraction
15
Tag to Molecule
41
Markush to Molecule
9
Organic Materials
OLED Property Extraction
13
Chemical Reaction
Drug Discovery
Reaction QA
15
Organic Materials
Reaction Mechanism QA
22
@@ 3.1 Table
To assess the table understanding capabilities of our model, Uni-SMART, compared to other leading models, we designed a diverse set of tasks across different domains. These tasks were specifically tailored to evaluate how well each model could interpret and extract information from tables, with the results presented in Table 2.
In our thorough assessment, the Uni-SMART showcased its exceptional proficiency in understanding and extracting table data from the scientific literature. Among six assessed tasks, it surpassed other models in four and delivered a competitive performance in the remainder. Notably, in the "Electrolyte Table QA" and "Polymer Property Extraction" tasks, Uni-SMART achieved Value Recall scores of 0.674 and 0.869, respectively, significantly outperforming its counterparts. These results highlight Uni-SMART’s exceptional skill in processing table data. However, Uni-SMART’s performance in the "Composition Extraction" task for Alloy Materials showed no significant difference compared to other models, only slightly leading. To further enhance its effectiveness, a possible approach could be to strengthen the model’s ability to learn domain-specific knowledge. Moreover, all models under-performed in the "Affinity Data Extraction" task within Drug Discovery and the "Electrolyte Solubility data extraction" task in Organic Materials, indicating the high complexity of these tasks. Such challenges might be attributed to the diversity of table formats and the presence of complex nested structures, such as tables embedding molecular structures. To navigate these difficulties, future directions could involve the development of more sophisticated table data comprehension technologies and the refinement of data preprocessing methods.
Table 2: Evaluation on table data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Alloy Materials
Composition Extraction
Value Recall
0.451
0.420
0.437
0.443
Drug Discovery
Affinity Data Extraction
Value Recall
0.320
0.274
0.359
0.175
Organic
Materials
Electrolyte Solubility data extraction
Value Recall
0.282
0.365
0.393
0.390
Electrolyte Table  QA
Value Recall
0.674
0.163
0.140
0.233
Polymer Property Extraction
Value Recall
0.869
0.762
0.286
0.506
OLED Property Extraction
Value Recall
0.526
0.520
0.479
0.394
@@ 3.2 Chart
Analysis of the results presented in Table 3 indicates that Uni-SMART significantly outperforms existing models like GPT-4, GPT-3.5, and Gemini in ChartQA tasks across diverse scientific domains, particularly excelling in the areas of alloy materials and organic materials, with leading scores of 0.667 and 0.733, respectively. However, despite Uni-SMART’s leading position, its advantage narrows in the domains of drug discovery and biology, where all models struggle to achieve high performance. This limitation is likely due to the complexity and specificity of charts in these fields, which demand higher levels of understanding and adaptability from the models. To overcome these challenges, future work should focus on augmenting the training dataset with a wider and more sophisticated assortment of chart data, aiming to enhance the model’s analytical depth and accuracy. Furthermore, exploring refined model architectures and specialized training methodologies, specifically designed to meet the complex requirements of chart analysis, is crucial for significantly elevating Uni-SMART’s performance in such demanding tasks.
Table 3: Evaluation on chart data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Alloy Materials
Alloy ChartQA
Accuracy
0.667
0.400
0.267
0.333
Biology
Biology ChartQA
Accuracy
0.467
0.333
0.400
0.400
Drug Discovery
Drug ChartQA
Accuracy
0.467
0.400
0.333
0.400
Organic Materials
Polymer ChartQA
Accuracy
0.733
0.667
0.333
0.007
3.3 Molecular Structure
@@
In scientific literature, molecular structures hold fundamental importance, especially within fields such as chemistry, pharmacology, and biology. The ability to comprehend molecular structures is crucial for analyzing research outcomes, predicting chemical behaviours, and innovating novel compounds [14, 15]. Deciphering the molecular structure of a newly discovered drug compound, for instance, can provide essential insights into its therapeutic potential and biological interactions, which is crucial for enhancing our understanding of related scientific literature.
@@
To evaluate our model’s ability to understand molecular structures, we carried out a series of tasks involving molecules, polymers, and Markush structures, common in chemistry and medicine studies. These tasks were designed to assess the model’s capability in deciphering information from representations of molecular structures. The execution of these tasks and their respective results are detailed in Table 4.
The Uni-SMART model showed outstanding performance in molecular structure understanding tasks, particularly in processing complex molecular structures within the scientific literature. This is evident from the results presented in Table 4. For example, in the "Tag to Molecule" task, Uni-SMART achieved an average similarity score of 0.275, significantly outperforming other models which struggled with this task (GPT-4 and Gemini achieved a score of 0, while GPT-3.5 only reached 0.007). In the "Markush to Molecule" task, its average similarity score of 0.629 was also markedly higher than its competitors. These results clearly indicate Uni-SMART’s advantages in understanding molecular structures in scientific literature.
However, despite its overall excellent performance across most tasks, Uni-SMART was slightly inferior to GPT-3.5 in the "Affinity Data Extraction" task. This task involves extracting affinity tables from scientific literature, requiring not only an understanding of complex, domain-specific language but also a deep analysis to match text with molecular structures and other modalities. This suggests that while Uni-SMART has strong capabilities in molecular structure understanding, there is still room for improvement in handling tasks requiring detailed information retrieval and cross-modal matching, indicating potential areas for further optimization.
Table 4: Evaluation on molecule data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Drug
Discovery
Affinity Data Extraction
Value Recall
0.320
0.274
0.359
0.175
Tag to Molecule
Mean Similarity
0.275
0.000
0.007
0.000
Molecule in Document
Accuracy
0.889
0.022
0.489
0.500
Markush to Molecule
Mean Similarity
0.629
0.545
0.459
0.199
Organic
Materials
OLED Property Extraction
Value Recall
0.526
0.520
0.479
0.394
@@ 3.4 Chemical Reactions
Understanding chemical reactions is crucial in the realm of scientific literature, particularly within chemistry and its related disciplines [16]. The ability to accurately parse and analyze the details of chemical reactions enhances readers’ grasp of complex material, enabling a deeper understanding of experimental results and theoretical discussions.
To assess Uni-SMART’s capability to comprehend chemical reactions, we designed two tasks aimed at challenging the model’s proficiency in interpreting chemical reactions. These tasks focus on understanding the reactants, products, and conditions of chemical reactions, as well as grasping the underlying mechanisms and clarifying the significance of these reactions within a broader scientific context. The results of these tasks, and a comparative analysis of Uni-SMART’s performance against other leading models, are summarized in Table 5.
Uni-SMART exhibited outstanding performance in "Reaction QA" within the drug discovery domain and "Reaction Mechanism QA" within the organic materials domain, achieving an accuracy of 0.4 and 0.445, respectively. These results significantly surpass those of GPT-4, GPT-3.5, and Gemini models. Such achievements highlight Uni-SMART’s considerable advantage in parsing chemical reactions, potentially attributed to its access to richer, more specialized training data sources, effective preprocessing of chemical expressions, and architecture designed specifically for multimodal data. Despite these strengths, there remains room for improvement in Uni-SMART’s comprehension of chemical reactions. Future enhancements may include strengthening the model’s understanding of reaction details and expanding the training dataset to encompass a broader range of reaction types and mechanisms. These efforts are expected to further augment Uni-SMART’s applicability in reading and interpreting scientific literature.
Table 5: Evaluation on chemical reactions data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Drug Discovery
Reaction QA
Accuracy
0.400
0.200
0.000
0.133
Organic Materials
Reaction Mechanism QA
Accuracy
0.445
0.382
0.345
0.373

### v2

@@ Abstract
In scientific research and its application, scientific literature analysis is crucial as it allows researchers to build on the work of others. However, the fast growth of scientific knowledge has led to a massive increase in scholarly articles, making in-depth literature analysis increasingly challenging and time-consuming. The emergence of Large Language Models (LLMs) has offered a new way to address this challenge. Known for their strong abilities in summarizing texts, LLMs are seen as a potential tool to improve the analysis of scientific literature. However, existing LLMs have their own limits. Scientific literature often includes a wide range of multimodal elements, such as tables, charts, and molecule, which are hard for text-focused LLMs to understand and analyze. This issue points to the urgent need for new solutions that can fully understand and analyze multimodal content in scientific literature. To answer this demand, we present Uni-SMART (Universal Science Multimodal Analysis and Research Transformer), an innovative model designed for in-depth understanding of multimodal scientific literature. Through rigorous quantitative evaluation across several domains, Uni-SMART demonstrates superior performance over other text-focused LLMs. Furthermore, our exploration extends to practical applications, including patent infringement detection and nuanced analysis of charts. These applications not only highlight Uni-SMART’s adaptability but also its potential to revolutionize how we interact with scientific literature.
@@ 1 Introduction
To rigorously assess the multimodal capabilities of Uni-SMART, a comparative analysis was conducted against several LLMs, such as GPT-4o, Gemini, and Claude3. Our assessment targets several data types critical to the comprehension of scientific documents: tables, charts, molecular structures, and chemical reactions. The results demonstrate Uni-SMART’s superior performance in all tested areas, especially in understanding and analyzing complex multimodal contents, thus highlighting its potential as a helpful assistant for scientific literature analysis.
In the following sections, Section 2 details Uni-SMART’s data sources and iterative training approach. Subsequently, Section 3 – Evaluation, presents detailed comparisons of Uni-SMART with several LLMs across a variety of modalities, showcasing its advanced capabilities in multimodal data interpretation. Then, Section 4 – Application showcases some specific applications of Uni-SMART in the scientific domain. Finally, Section 5 discusses Uni-SMART’s limitations and future research avenues, along with its potential impact on scientific research and technological advancement.
@@ 2 Method
  • Multimodal Learning: During the initialization phase, the parsing model is trained with a limited set of multimodal data to recognize and extract diverse information elements from scientific literature. The output is formatted in a custom text format, similar to html format, to effectively represent multimodal elements.
  • LLM SFT: A series of valuable queries are constructed, particularly in multimodal scenarios. With Multimodal Retrieval-Augmented Generation (RAG), relevant content is recalled from the literature based on queries. Answers are built based on the queries and retrieved content. The query-answer data is then used to fine-tune the LLM. This process helps the LLM adapt to the custom input format and improves its ability to follow instructions.
  • User Feedback: The parsing model and SFT-enhanced LLM are deployed in real applications, facilitating the collection of user feedback. Samples receiving positive feedback are subsequently filtered and incorporated into the data enhancement, while those with negative feedback are subject to expert annotation before being integrated into the data enhancement process.
  • Expert Annotation: Samples with negative feedback are carefully annotated by human experts. This step ensures that these models learns from their mistakes, with semi-automated tools assisting in this process to enhance efficiency. Negative feedback cases typically fall into the following categories:
    1. 1.
      Multimodal element recognition errors;
    2. 2.
      Recall content errors;
    3. 3.
      Poor instruction-following.
    Detailed analysis of these error types facilitates targeted improvements.
  • Data Enhancement: Finally, the annotated data, along with partial samples that received positive feedback, are added to the training dataset for data enhancement. The pipeline is optimized based on different types of negative feedback:
    1. 1.
      Multimodal element recognition errors: expand the training data for the parsing model;
    2. 2.
      Recall content errors: optimize the Multimodal RAG scheme;
    3. 3.
      Poor instruction-following: enlarge and refine the dataset used for LLM SFT.
This iterative process is repeated to continually optimize Uni-SMART’s overall performance. And significantly enhances Uni-SMART’s performance in a variety of challenging tasks, such as information extraction, complex element identification, scientific literature understanding/analysis, and multimodal understanding/reasoning.
@@ 3 Evaluation
In this section, we perform a detailed quantitative evaluation on the capabilities of Uni-SMART and various available LLMs across modalities. Table 1 presents the statistics of the benchmarks. For details of the specific evaluation tasks, please refer to SciAssess [10].
For our experiments, different methods are employed depending on the source of the models. Closed-source models are accessed via API calls, while open-source models are obtained from Hugging Face, deployed, and tested. Evaluation tasks that require an article as context involve converting the PDF content to text for input into the LLMs. If an LLM offers a built-in PDF parsing interface, we utilize this feature; otherwise, PyPDF2 is used for PDF parsing. Notably, Uni-SMART can directly read PDF files, allowing us to upload the original documents and pose questions directly to the model. Due to input length limitations of the LLMs, tasks that require the full text of an article as context are executed in a zero-shot manner, whereas tasks that do not require such extensive context are tested using few-shot approaches.
The final results of these experiments are systematically presented in Table 2, illustrating the performance of each model across the evaluated tasks.
Table 1: Statistics of evaluation datasets
Modality Domain Task Question Type Metric # Question
Table Alloy Materials Composition Extraction Table Extraction Table Recall 244
Drug Discovery Affinity Extraction Table Extraction Table Recall 40
Organic Materials Electrolyte Table QA Multi Choices Accuracy 100
OLED Property Extraction Table Extraction Table Recall 13
Polymer Property Extraction Table Extraction Table Recall 109
Solubility Extraction Table Extraction Table Recall 100
Chart Alloy Materials Alloy Chart QA Multi Choices Accuracy 15
Biology Biology Chart QA Multi Choices Accuracy 99
Drug Discovery Drug Chart QA Multi Choices Accuracy 15
Organic Materials Polymer Chart QA Multi Choices Accuracy 15
Molecule Drug Discovery Affinity Data Extraction Table Extraction Table Recall 40
Tag to Molecule Molecule Generation Molecule Similarity 50
Markush to Molecule Molecule Generation Molecule Similarity 37
Molecule in Document YES/NO Accuracy 50
Organic Materials OLED Property Extraction Table Extraction Table Recall 13
Reaction Drug Discovery Reaction QA Multi Choices Accuracy 95
Organic Materials Reaction Mechanism QA Multi Choices Accuracy 22
Table 2: Performance Comparison of LLMs on Diverse Tasks.
Modality Task Uni-SMART GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3 DeepSeek Qwen2 Command $\text{R}^{{\text{+}}}$
Table Composition Extraction 0.511 0.484 0.458 0.112 0.127 0.495 0.304 0.239 0.212 0.389 0.423 0.128
Affinity Extraction 0.200 0.072 0.042 0.025 0.040 0.097 0.050 0.040 0.064 0.017 0.075 0.043
Electrolyte Table QA 0.850 0.890 0.790 0.370 0.670 0.870 0.710 0.880 0.460 0.720 0.620 0.450
OLED Property Extraction 0.490 0.336 0.406 0.201 0.037 0.477 0.259 0.093 0.263 0.292 0.392 0.234
Polymer Property Extraction 0.705 0.692 0.681 0.329 0.705 0.629 0.514 0.606 0.536 0.652 0.636 0.171
Solubility Extraction 0.468 0.435 0.440 0.410 0.363 0.426 0.371 0.397 0.399 0.432 0.400 0.351
Chart Alloy Chart QA 0.933 0.533 0.600 0.333 0.333 0.400 0.467 0.667 0.467 0.333 0.400 0.200
Biology Chart QA 0.616 0.580 0.480 0.390 0.545 0.505 0.480 0.616 0.520 0.545 0.515 0.535
Drug Chart QA 0.600 0.333 0.400 0.067 0.400 0.200 0.533 0.533 0.400 0.400 0.400 0.533
Polymer Chart QA 0.933 0.800 0.667 0.400 0.800 0.467 0.867 0.800 0.867 0.733 0.933 0.800
Molecule Affinity Extraction 0.200 0.072 0.042 0.025 0.040 0.097 0.050 0.040 0.064 0.017 0.075 0.043
Tag to Molecule 0.188 0.040 0.022 0.000 0.016 0.035 0.094 0.169 0.034 0.014 0.000 0.031
Markush to Molecule 0.686 0.642 0.654 0.431 0.504 0.675 0.239 0.526 0.491 0.470 0.379 0.376
Molecule in Document 0.720 0.580 0.700 0.500 0.460 0.480 0.560 0.640 0.680 0.460 0.460 0.460
OLED Property Extraction 0.490 0.336 0.406 0.201 0.037 0.477 0.259 0.093 0.263 0.292 0.392 0.234
Reaction Reaction QA 0.768 0.705 0.674 0.442 0.253 0.663 0.442 0.305 0.611 0.368 0.442 0.316
Reaction Mechanism QA 0.682 0.545 0.636 0.455 0.545 0.455 0.636 0.727 0.500 0.545 0.591 0.591
@@ 3.1 Table
To assess the table understanding capabilities of our model, Uni-SMART, compared to other LLMs, we designed a diverse set of tasks across different domains. These tasks were specifically tailored to evaluate how well each model could interpret and extract information from tables.
In our thorough assessment, the Uni-SMART showcased its exceptional proficiency in understanding and extracting table data from the scientific literature. Among six assessed tasks, it surpassed other models in five and delivered a competitive performance in the remainder. Notably, in the "Composition Extraction" and "Solubility Extraction" tasks, Uni-SMART achieved "Table Recall" of 0.511 and 0.468, respectively, significantly outperforming its counterparts. The objectives of these two tasks are to extract necessary information from tables within articles and organize it into a specified format. These results indicate that Uni-SMART excels in handling and understanding tabular data, particularly in information extraction and formatting tasks.
However, in the Electrolyte Table QA task, Uni-SMART’s performance was slightly below the state-of-the-art model (GPT-4o). In this task, LLMs are required to answer questions about details in the table with the question type of multiple-choice, indicating that Uni-SMART still has room for improvement in understanding the details within tables.
Moreover, all models under-performed in the "Affinity Extraction" task, indicating the high complexity of these tasks. This task requires LLMs to organize molecules and corresponding experimental result from articles. It necessitates the model’s ability to integrate molecules with table information and demands strong long-context understanding capabilities. To better address these issues in the future, a possible direction is to enhance LLMs’ abilities in long context comprehension and information matching.
@@ 3.2 Chart
Analysis of the results presented in Figure 3 indicates that Uni-SMART significantly outperforms existing models in Chart QA tasks across diverse scientific domains, particularly excelling in the tasks of Alloy Chart QA and Drug Chart QA, with leading scores of 0.933 and 0.600, respectively. However, both in the Biology Chart QA, where the Gemini model scored 0.616, and in the Polymer Chart QA task, where Qwen2 also achieved a top score of 0.933—matching Uni-SMART’s performance in both cases—it is evident that these models have strong capabilities in analyzing scientific charts. These results underscore the intense competition among different models in the context of chart understanding. The future direction for development likely focuses on enhancing accuracy and applicability across various scientific domains.
3.3 Molecule
@@
In scientific literature, molecule or molecular structures hold fundamental importance, especially within fields such as chemistry, pharmacology, and biology. Comprehending molecular structures is crucial for analyzing research outcomes, predicting chemical behaviors, and innovating novel compounds [14, 15]. Deciphering the molecular structure of a newly discovered drug compound, for instance, can provide essential insights into its therapeutic potential and biological interactions, which is crucial for enhancing our understanding of related scientific literature.
@@
To evaluate our model’s ability to understand molecular structures, we carried out a series of tasks involving molecules, polymers, and Markush structures, which are common in chemistry and medicine studies. These tasks were designed to assess the model’s capability in deciphering information from representations of molecular structures.
The Uni-SMART model showed outstanding performance in molecular structure understanding tasks, particularly in processing complex molecular structures within the scientific literature. For example, in the "Molecule in Document" task, Uni-SMART achieved an accuracy of 0.720, significantly outperforming other models. This task requires LLMs to determine whether a molecular structure has appeared in the literature, demonstrating Uni-SMART’s strong capabilities in accurate molecule identification and information retrieval.
Such achievements highlight Uni-SMART’s considerable advantage in parsing molecular structure, potentially attributed to its access to richer, more specialized training data sources, effective preprocessing of molecular structure, and architecture designed specifically for multimodal data. Despite these strengths, there remains room for improvement in Uni-SMART’s comprehension of molecular structure. Future enhancements may include strengthening the model’s understanding of molecule details and expanding the training dataset to encompass a broader range of molecule types. These efforts are expected to further augment Uni-SMART’s applicability in reading and interpreting scientific literature, especially those involving molecular structure.
@@ 3.4 Chemical Reactions
Understanding chemical reactions is crucial in scientific literature, particularly within chemistry and its related disciplines [16]. The ability to accurately parse and analyze the details of chemical reactions enhances readers’ grasp of complex material, enabling a deeper understanding of experimental results and theoretical discussions.
To assess Uni-SMART’s capability to comprehend chemical reactions, we designed two tasks aimed at challenging the model’s proficiency in interpreting chemical reactions. These tasks focus on understanding the reactants, products, and conditions of chemical reactions, as well as grasping the underlying mechanisms and clarifying the significance of these reactions within a broader scientific context.
Uni-SMART exhibited outstanding performance in "Reaction QA" task, achieving an accuracy of 0.768. However, in the "Reaction Mechanism QA" task, Uni-SMART’s performance was slightly below that of the Gemini. This task requires LLMs to accurately parse and interpret the step-by-step mechanisms of chemical reactions from the article, indicating that Uni-SMART still has room for improvement in understanding complex reactions.
Figure 3: Comparative Performance of LLMs Across Different Modalities.